Discuss energy stored in a capacitance
1
Expert's answer
2019-11-12T17:26:57-0500

Capacitors store electric energy. The main expression for the energy stored is


"E=\frac{1}{2}CV^2."

Since


"C=\frac{Q}{V},"

the energy can be also expressed as


"E=\frac{1}{2}QV=\frac{1}{2}\frac{Q^2}{C}."

Capacitors can quickly consume electric energy (or charge) and give it out, and this can be considered as a key difference between capacitors and chemical sources of electrical energy.